how many cubes with two inch edges will be needed to fill a rectangular prism with dimension of 2 feet by 3 feet by 4 feet
Answer by checkley77(12844)

A 2 in. cube has 2^3=8 in^3 volume.
Your rectangular prism has 2*3*4=24 ft^3 volume
1 ft^3=1,728 in^3.
24*1,728/8=41,472/8=5,184 2 in. cubes will fit in the prism.